solve the system of three variable linear equations

x + 2y = 1

3x + 2y + 4z = 7

-2x + y - 2z = -1



1
Expert's answer
2021-03-29T13:54:20-0400

The augmented matrix of the given system is:


"\begin{bmatrix}
 1 & 2 &0:1\\
 3 & 2&4:7\\
 -2&1&-2:-1
\end{bmatrix}"

Now, "R_3\rightarrow3R_1-R_3;\ \ R_3\rightarrow2R_1+R_3"

"\begin{bmatrix}
 1 & 2 &0:1\\
 0 & 4&-4:4\\
 0&5&-2:1
\end{bmatrix}"

"R_2\rightarrow R_2/4;"

"\begin{bmatrix}
 1 & 2 &0:1\\
 0 & 1&-1:1\\
 0&5&-2:1
\end{bmatrix}"

"R_1\rightarrow R_1-2R_2;\ \ R_3\rightarrow R_3-5R_2"

"\Rightarrow \begin{bmatrix}
 1 & 0 &2:3\\
 0 & 1&-1:1\\
 0&0&3:6
\end{bmatrix}"

Comparing on third row,

"3z=6\Rightarrow [z=2]"

On second row,


"y-z=-1\Rightarrow y=-1+z\\ \Rightarrow [y=1]"

On first row,


"x=3-2z\Rightarrow [x=-1]"

"\Rightarrow\begin{bmatrix}
 x \\
 y\\
z
\end{bmatrix}=\begin{bmatrix}
 -1\\
 1\\
2
\end{bmatrix}"